# On the Convergence of HalpernSGD

### v1

@@ 4 Convergence Analysis of HalpernSGD
$\lim_{k\to 0}\Bigl(2\langle u-x^{*},\,\mathbb{E}[x_{n_{k}}-x^{*}]\rangle-\beta_{n_{k}}\,\mathbb{E}\|\nabla f(x_{n_{k}})\|^{2}\Bigr)=0.$ (19)
Observe that Lemma 4.1 implies that $2\langle u-x^{*},\,\mathbb{E}[x_{n_{k}}-x^{*}]\rangle$ remains bounded, while $\beta_{n_{k}}\sim\epsilon_{n_{k}}/\alpha_{n_{k}}$ diverges.
These facts readily implies that $\mathbb{E}\|\nabla f(x_{n_{k}})\|^{2}\to 0$ and, by eventually passing to further subsequences, that
$\nabla f(x_{n_{k}})\to 0\text{ a.s.}$ (20)
as well.
Now, observe that for every $n$,
$\mathbb{E}\bigl[\|x_{n+1}-x_{n}\|^{2}\,\big|\,\mathcal{F}_{n}\bigr]=\mathbb{E}\bigl[\bigl\|\alpha_{n}(u-x_{n})+(1-\alpha_{n})\bigl(-\epsilon_{n}(\nabla f(x_{n})+U_{n}\bigr))\bigr\|^{2}\big|\,\mathcal{F}_{n}\bigr].$
Using the inequality $\|a+b\|^{2}\leq 2\|a\|^{2}+2\|b\|^{2}$, we obtain
$\mathbb{E}\bigl[\|x_{n+1}-x_{n}\|^{2}\,\big|\,\mathcal{F}_{n}\bigr]\;\leq\;2\alpha_{n}^{2}\|u-x_{n}\|^{2}+2\epsilon_{n}^{2}\bigl(\|\nabla f(x_{n})\|^{2}+\sigma^{2}\bigr).$
Since $\mathbb{E}\|u-x_{n}\|^{2}$ and $\mathbb{E}\|\nabla f(x_{n})\|^{2}$ are bounded, by Lemma 4.1, while $\sum_{n}\alpha_{n}^{2}<\infty$ and $\sum_{n}\epsilon_{n}^{2}<\infty$, taking total expectations we deduce that
$\sum_{n=0}^{\infty}\mathbb{E}\|x_{n+1}-x_{n}\|^{2}<\infty,$
Hence, by a standard convergence argument (see, e.g., [15]), it follows that
$\|x_{n+1}-x_{n}\|\;\to\;0\quad\text{almost surely.}$
Moreover, since $\nabla f$ is $L$-Lipschitz, $\|\nabla f(x_{n+1})-\nabla f(x_{n})\|\;\leq\;L\,\|x_{n+1}-x_{n}\|,$ it holds as well
$\|\nabla f(x_{n+1})-\nabla f(x_{n})\|\to 0\text{ a.s.}$
Returning to (20) we can finally conclude that
$\|\nabla f(x_{n})\|\;\to\;0\quad\text{almost surely.}$
Let $(x_{n_{k}})$ be a subsequence such that
[…]
which exists and is finite since the sequence $\bigl(\mathbb{E}\langle x_{n}-x^{*},\,u-x^{*}\rangle\bigr)$ is bounded.
By passing, if necessary, to a further subsequence, we may assume that for almost every $\omega\in\Omega$,
$\langle x_{n_{k}}(\omega)-x^{*},\,u-x^{*}\rangle\;\longrightarrow\;\langle P(\omega)-x^{*},\,u-x^{*}\rangle$
for some cluster point $P(\omega)$ of $(x_{n}(\omega))$.
Since $\nabla f(x_{n_{k}})\to 0$ almost surely, it follows that each limit point $P(\omega)$ satisfies $\nabla f(P(\omega))=0$; that is, $P(\omega)$ is a critical point of $f$. Moreover, since
$\sup_{n}\mathbb{E}\|x_{n}-x^{*}\|^{2}\|u-x^{*}\|^{2}<\infty,$
the sequence of random variables
$X_{n}:=\langle x_{n}-x^{*},\,u-x^{*}\rangle$
is bounded in $L^{2}$. By the de la Vallée–Poussin criterion, this implies that $\{X_{n}\}$ is uniformly integrable Let $(x_{n_{k}})$ be a subsequence such that
[…]
for some measurable random variable $P$. Then, for those $\omega$,
$X_{n_{k}}(\omega)=\langle x_{n_{k}}(\omega)-x^{*},\,u-x^{*}\rangle\longrightarrow\langle P(\omega)-x^{*},\,u-x^{*}\rangle.$
This last, together with the already proved uniform integrability implies the convergence in $L^{1}$ by means of Vitali’s Theorem. Hence, we deduce that
$\lim_{k\to\infty}\mathbb{E}\,\langle x_{n_{k}}-x^{*},\,u-x^{*}\rangle=\mathbb{E}\,\langle P-x^{*},\,u-x^{*}\rangle.$
Since $\nabla f(P(\omega))=0,$ $P(\omega)\in S$ for almost every $\omega,$ and since $x^{*}=P_{S}(u)$, it follows that
$\langle P(\omega)-x^{*},\,u-x^{*}\rangle\leq 0\quad\text{for almost every }\omega,$
and therefore
$\limsup_{n\to\infty}\mathbb{E}\,\langle x_{n}-x^{*},\,u-x^{*}\rangle\;\leq\;0.$
Returning to (17), we rewrite it as
$\mathbb{E}\|x_{n+1}-x^{*}\|^{2}$ $\leq(1-\alpha_{n})\mathbb{E}\|x_{n}-x^{*}\|^{2}+\alpha_{n}^{2}\|u-x^{*}\|^{2}\quad+\epsilon_{n}^{2}\sigma^{2}$
@@
$+2\alpha_{n}\langle u-x^{*},\,\mathbb{E}[x_{n}-x^{*}]\rangle+2\alpha_{n}\epsilon_{n}L\sqrt{C}\,\|u-x^{*}\|.$
We observe that the sequences $(\alpha_{n}^{2})$, $(\epsilon_{n}^{2})$, and $(\alpha_{n}\epsilon_{n})$ are summable, while $\limsup_{n\to\infty}\langle u-x^{*},\,\mathbb{E}[x_{n}-x^{*}]\rangle\leq 0.$ Hence, by applying the Xu Lemma (see [35, Lemma 2.1]), we conclude that
@@
Consequently, there exists a subsequence $(x_{n_{k}})$ such that $\|x_{n_{k}}-x^{*}\|\to 0$ almost surely. Since we have already established that $\|x_{n+1}-x_{n}\|\to 0$ almost surely, this asymptotic regularity implies that the entire sequence converges:
$x_{n}\;\longrightarrow\;x^{*}\quad\text{almost surely.}$
@@ 5 Conclusion and future perspectives
This publication was partially funded by the PhD program in Mathematics and Computer Science at University of Calabria, Cycle XXXVIII with the support of a scholarship financed by DM 351/2022 (CUP H23C22000440007), based on the NRPP funded by the European Union.

### v2

@@ 4 Convergence Analysis of HalpernSGD
$\kappa_{n}:=\frac{(1-\alpha_{n})^{2}}{\alpha_{n}}\Bigl(\frac{2\epsilon_{n}}{L}-\epsilon_{n}^{2}\Bigr)\geq 0,\qquad b_{n}:=2(1-\alpha_{n})\langle u-x^{*},\,\mathbb{E}[x_{n}-x^{*}]\rangle-\kappa_{n}\,\mathbb{E}\|\nabla f(x_{n})\|^{2},$
$c_{n}:=\alpha_{n}^{2}\|u-x^{*}\|^{2}+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\sigma^{2}+2\alpha_{n}\epsilon_{n}(1-\alpha_{n})L\sqrt{C}\,\|u-x^{*}\|.$
Then (16) can be rewritten as
$a_{n+1}\leq(1-\alpha_{n})a_{n}+\alpha_{n}b_{n}+c_{n}.$ (17)
By the step-size assumptions, $\sum_{n}\alpha_{n}^{2}<\infty$, $\sum_{n}\epsilon_{n}^{2}<\infty$ and $\sum_{n}\alpha_{n}\epsilon_{n}<\infty$, hence $\sum_{n}c_{n}<\infty$.
Step 1: $\limsup_{n\to\infty}b_{n}\leq 0$. First note that, by Lemma 4.1, there exists $C_{0}>0$ such that $\sup_{n}a_{n}\leq C_{0}$. Hence, by Cauchy–Schwarz,
$\big|2(1-\alpha_{n})\langle u-x^{*},\,\mathbb{E}[x_{n}-x^{*}]\rangle\big|\leq 2\|u-x^{*}\|\,\sqrt{a_{n}}\leq 2\|u-x^{*}\|\sqrt{C_{0}}=:M.$ (18)
Moreover, by the coupling $(1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\leq 1$ we have $\epsilon_{n}^{2}L^{2}\leq\alpha_{n}<1$, hence $\epsilon_{n}\leq 1/L$ for all $n$. Therefore
$\Bigl(\frac{2\epsilon_{n}}{L}-\epsilon_{n}^{2}\Bigr)=\epsilon_{n}\Bigl(\frac{2}{L}-\epsilon_{n}\Bigr)\geq\frac{\epsilon_{n}}{L}.$
Therefore,
$\kappa_{n}\;\geq\;\frac{(1-\alpha_{n})^{2}}{\alpha_{n}}\cdot\frac{\epsilon_{n}}{L}\;\xrightarrow[n\to\infty]{}\;+\infty\qquad\text{since }\frac{\epsilon_{n}}{\alpha_{n}}\to+\infty.$ (19)
We claim that $\limsup_{n\to\infty}b_{n}\leq 0$. Suppose by contradiction that $\limsup_{n\to\infty}b_{n}=\delta>0$. Then there exists a subsequence $(n_{k})$ such that $b_{n_{k}}\to\delta$ and hence $b_{n_{k}}\geq\delta/2$ for $k$ large enough. Using (18),
@@
$\kappa_{n_{k}}\,\mathbb{E}\|\nabla f(x_{n_{k}})\|^{2}=2(1-\alpha_{n_{k}})\langle u-x^{*},\mathbb{E}[x_{n_{k}}-x^{*}]\rangle-b_{n_{k}}\leq M-\frac{\delta}{2},$
so, by (19),
$\mathbb{E}\|\nabla f(x_{n_{k}})\|^{2}\longrightarrow 0.$ (20)
In particular, by Markov’s inequality, $\nabla f(x_{n_{k}})\to 0$ in probability.
Since $\sup_{n}\mathbb{E}\|x_{n}-x^{*}\|^{2}<\infty$, the family $\{x_{n}\}$ is tight.
Thus, passing to a further subsequence (not relabeled), we may assume that the laws of $x_{n_{k}}$ converge weakly to some probability measure $\mu$ on $\mathbb{R}^{d}$. Equivalently, there exists an $\mathbb{R}^{d}$-valued random vector $P$ (defined on some probability space) with law $\mu$ such that
$x_{n_{k}}\Rightarrow P\qquad\text{in distribution.}$
By the continuous mapping theorem, $\nabla f(x_{n_{k}})\Rightarrow\nabla f(P)$. On the other hand, $\nabla f(x_{n_{k}})\to 0$ in probability implies $\nabla f(x_{n_{k}})\Rightarrow 0$, hence $\nabla f(P)=0$ almost surely, i.e. $P\in S$ almost surely.
Now define $X_{n}:=\langle x_{n}-x^{*},\,u-x^{*}\rangle$. By Cauchy–Schwarz, $\sup_{n}\mathbb{E}|X_{n}|^{2}\leq\|u-x^{*}\|^{2}\sup_{n}a_{n}<\infty$, hence $\{X_{n}\}$ is uniformly integrable.
Since $x_{n_{k}}\Rightarrow P$, we have $X_{n_{k}}\Rightarrow X:=\langle P-x^{*},u-x^{*}\rangle$. Since $\{X_{n_{k}}\}$ is uniformly integrable and $X_{n_{k}}\Rightarrow X$, we have
$\mathbb{E}X_{n_{k}}\longrightarrow\mathbb{E}X.$
Since $P(\omega)\in S$ a.s. and $x^{*}=P_{S}(u)$, the projection characterization gives $\langle u-x^{*},P-x^{*}\rangle\leq 0$ a.s., hence $\mathbb{E}\langle P-x^{*},u-x^{*}\rangle\leq 0$. Therefore,
$\limsup_{k\to\infty}2(1-\alpha_{n_{k}})\langle u-x^{*},\mathbb{E}[x_{n_{k}}-x^{*}]\rangle\leq 0.$
Together with (20) and $\kappa_{n_{k}}\geq 0$, this implies
$\limsup_{k\to\infty}b_{n_{k}}\leq 0,$
contradicting $b_{n_{k}}\to\delta>0$. Hence $\limsup_{n\to\infty}b_{n}\leq 0$.
Step 2: apply Xu’s lemma. We have shown that (17) holds with $\sum_{n}c_{n}<\infty$ and $\limsup b_{n}\leq 0$. Since also $\sum_{n}\alpha_{n}=\infty$ and $\alpha_{n}\in(0,1)$, Xu’s lemma (see [43, Lemma 2.1]) yields $a_{n}\to 0$, i.e.
@@
Convergence in probability follows from Chebyshev’s inequality. Finally, choose integers $n_{k}\uparrow\infty$ such that $\mathbb{E}\|x_{n_{k}}-x^{*}\|^{2}\leq 2^{-k}$. Then, for any $\nu>0$, by Markov,
$\sum_{k=1}^{\infty}\mathbb{P}\big(\|x_{n_{k}}-x^{*}\|>\nu\big)\leq\frac{1}{\nu^{2}}\sum_{k=1}^{\infty}\mathbb{E}\|x_{n_{k}}-x^{*}\|^{2}\leq\frac{1}{\nu^{2}}\sum_{k=1}^{\infty}2^{-k}<\infty.$
By Borel–Cantelli, $\|x_{n_{k}}-x^{*}\|\to 0$ almost surely. ∎
The following lemma is a standard consequence of Baillon–Haddad cocoercivity; see, e.g., [3, 4].
Lemma 4.3 (Nonexpansiveness of a gradient step).
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be convex and $L$-smooth. Then for every $\epsilon\in[0,2/L]$, the mapping $T_{\epsilon}(x):=x-\epsilon\nabla f(x)$ is nonexpansive:
$\|T_{\epsilon}(x)-T_{\epsilon}(y)\|\leq\|x-y\|\qquad\forall x,y\in\mathbb{R}^{d}.$
Proof.
Using cocoercivity (11), for all $x,y$ we have
$\|T_{\epsilon}(x)-T_{\epsilon}(y)\|^{2}=\|x-y-\epsilon(\nabla f(x)-\nabla f(y))\|^{2}$
$\|x-y\|^{2}-2\epsilon\langle\nabla f(x)-\nabla f(y),x-y\rangle+\epsilon^{2}\|\nabla f(x)-\nabla f(y)\|^{2}$
$\leq\|x-y\|^{2}-\Bigl(\frac{2\epsilon}{L}-\epsilon^{2}\Bigr)\|\nabla f(x)-\nabla f(y)\|^{2}\leq\|x-y\|^{2},$
whenever $\epsilon\leq 2/L$. Taking square roots gives the claim. ∎
Theorem 4.4 (Almost sure convergence).
Assume the hypotheses of Theorem 4.2. Then
$x_{n}\longrightarrow x^{*}\qquad\text{almost surely.}$
Proof.
Step 0: define the noise-free reference sequence. Let $\{\bar{x}_{n}\}$ be the noise-free Halpern gradient sequence defined by
$\bar{x}_{n+1}:=\alpha_{n}u+(1-\alpha_{n})\bigl(\bar{x}_{n}-\epsilon_{n}\nabla f(\bar{x}_{n})\bigr),\qquad\bar{x}_{0}:=x_{0}.$
For each fixed initial point $\bar{x}_{0}\in\mathbb{R}^{d}$, the deterministic recursion above converges to $x^{*}=P_{S}(u)$ (this is the deterministic special case of Theorem 4.2, obtained by setting $U_{n}\equiv 0$; since the recursion is deterministic, the conclusion is pointwise). Hence, for the (possibly random) initial $x_{0}$, we still have
$\bar{x}_{n}(\omega)\longrightarrow x^{*}\qquad\text{for all }\omega\in\Omega.$ (21)
Step 1: the stochastic iterates track the noise-free iterates almost surely. Define the error $e_{n}:=x_{n}-\bar{x}_{n}$. Using the updates for $x_{n+1}$ and $\bar{x}_{n+1}$, we get
$e_{n+1}=(1-\alpha_{n})\Bigl((x_{n}-\epsilon_{n}\nabla f(x_{n}))-(\bar{x}_{n}-\epsilon_{n}\nabla f(\bar{x}_{n}))-\epsilon_{n}U_{n}\Bigr)=(1-\alpha_{n})\bigl(T_{\epsilon_{n}}(x_{n})-T_{\epsilon_{n}}(\bar{x}_{n})-\epsilon_{n}U_{n}\bigr),$
where $T_{\epsilon}(x)=x-\epsilon\nabla f(x)$.
Taking squared norms and conditional expectation with respect to $\mathcal{F}_{n}$, and using that $\mathbb{E}[U_{n}\mid\mathcal{F}_{n}]=0$, we obtain
$\mathbb{E}\bigl[\|e_{n+1}\|^{2}\mid\mathcal{F}_{n}\bigr]$ $=(1-\alpha_{n})^{2}\Bigl(\|T_{\epsilon_{n}}(x_{n})-T_{\epsilon_{n}}(\bar{x}_{n})\|^{2}+\epsilon_{n}^{2}\,\mathbb{E}[\|U_{n}\|^{2}\mid\mathcal{F}_{n}]\Bigr).$
Note that $(1-\alpha_{n})+\epsilon_{n}^{2}L^{2}\leq 1$ implies $\epsilon_{n}^{2}L^{2}\leq\alpha_{n}<1$, hence $\epsilon_{n}\leq 1/L<2/L$, so Lemma 4.3 applies to $T_{\epsilon_{n}}$. By Lemma 4.3 (nonexpansiveness of $T_{\epsilon_{n}}$) and by $\mathbb{E}[\|U_{n}\|^{2}\mid\mathcal{F}_{n}]\leq\sigma^{2}$, we get
$\mathbb{E}\bigl[\|e_{n+1}\|^{2}\mid\mathcal{F}_{n}\bigr]\leq(1-\alpha_{n})^{2}\|e_{n}\|^{2}+(1-\alpha_{n})^{2}\epsilon_{n}^{2}\sigma^{2}\leq(1-\alpha_{n})\|e_{n}\|^{2}+\epsilon_{n}^{2}\sigma^{2},$
where we used $(1-\alpha_{n})^{2}\leq(1-\alpha_{n})$ and $(1-\alpha_{n})^{2}\leq 1$.
Set $V_{n}:=\|e_{n}\|^{2}$. Then
$\mathbb{E}[V_{n+1}\mid\mathcal{F}_{n}]\leq V_{n}-\alpha_{n}V_{n}+\epsilon_{n}^{2}\sigma^{2}.$
Since $\sum_{n}\epsilon_{n}^{2}<\infty$, the Robbins–Siegmund almost-supermartingale theorem (see, e.g., [18]) implies that $V_{n}$ converges almost surely and
$\sum_{n=0}^{\infty}\alpha_{n}V_{n}<\infty\qquad\text{a.s.}$
Because $\sum_{n}\alpha_{n}=\infty$ and $V_{n}\geq 0$, the latter forces $\liminf_{n\to\infty}V_{n}=0$. Since $V_{n}$ also converges a.s., we conclude that
$\|e_{n}\|^{2}=V_{n}\longrightarrow 0\qquad\text{a.s.}$ (22)
Step 2: conclude. By (21) and (22),
$\|x_{n}-x^{*}\|\leq\|x_{n}-\bar{x}_{n}\|+\|\bar{x}_{n}-x^{*}\|\longrightarrow 0\qquad\text{a.s.}$
Hence $x_{n}\to x^{*}$ almost surely. ∎
@@ 5 Conclusion and future perspectives
This publication was partially funded by the PhD program in Mathematics and Computer Science at University of Calabria, Cycle XXXVIII with the support of a scholarship financed by DM 351/2022 (CUP H23C22000440007), based on the NRRP funded by the European Union.